One grocery clerk can stock a shelf in 35 min. A second clerk requires 40 min to stock the same shelf. How long would it take to stock the shelf if the two clerks worked together?
Answer by mananth(16946)

one clerk 35 minutes
other clerk 40 minutes
..
first clerk 1/35 of the job in 1 minute
second clerk 1/40 of the job in 1 minute
..
together they do 1/35 + 1/40 of the job in 1 minute
=(40+35)/40*35
=75/1400
=3/56 of the job in 1 minute
so they will take 56/3 = 18.66 hours together.
...
m.ananth@hotmail.ca